OpenStudy (anonymous):

Solve the separable differential equation for u : du/dt =e^(6u+6t) use the following initial condition u(0)=10

OpenStudy (anonymous):

e^(-6u) du=e^(6t)dt e^(-6u)/-6=e^(6t)+C puting intial condition e^(-60)/-6=1+C so C=-(1+1/6e^(60))

OpenStudy (anonymous):

but i need to solve to u not just c

OpenStudy (anonymous):

you can write like that e^(-6u)=-6e^(6t)-6C take anti log on both side -6u=ln(-6e^(6t)-6C) u=ln(1/(-6e^(6t)-6C)^6)

OpenStudy (anonymous):

is not right :(